# ToxSearch: Evolving Prompts for Toxicity Search in Large Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities for many tasks, but they also pose risks by potentially generating toxic content [^1] This paper contains disturbing language presented solely for safety evaluation.. Toxicity is defined as “a rude, disrespectful, or unreasonable comment likely to make someone leave a discussion” [ref_url1]. Even when safety-aligned, e.g., via fine-tuning or reinforcement learning from human feedback (RLHF), these models are not foolproof as adversarial inputs can still elicit toxic outputs. In particular, prompt-based attacks have emerged as a central concern because carefully crafted prompts can slip past safety filters. Identifying these prompts and building automated methods to discover them are essential [corbo2025toxic, Perez2022RedTL, Srivastava2023NoOT].
Growing concerns about LLM safety led to research on jailbreak attacks [corbo2025toxic, zou2023universal], but manual techniques are hard to scale and systematically evaluate. Zou et al. demonstrated that automatic prompt attacks are possible and found a universal adversarial suffix via greedy and gradient-based search that caused multiple LLMs to comply with harmful requests at high success rates [zou2023universal]. However, gradient-based approaches such as this usually need differentiable surrogates and often yield unnatural token sequences, limiting their usefulness for text.
@@
To overcome the limitations of manual and gradient-based attacks, researchers have turned to evolutionary algorithms (EAs), which frame prompt generation as an optimization task  [corbo2025toxic, Yu2023GPTFUZZERRT, Yang2023InstOptimaEM, Liu2023AutoDANGS]. A major benefit of EAs is that they require no access to model internals. For instance, EvoTox  [corbo2025toxic] casts toxicity elicitation as a $(1+\lambda)$ evolutionary strategy in which one LLM is the system under test and another LLM acts as a prompt generator that mutates a parent prompt to increase the system model’s toxicity. EvoTox was able to induce highly toxic outputs even from aligned models, outperforming random search and baseline jailbreak prompts on the AdvBench dataset  [zou2023universal]. Likewise, GPTFuzzer  [Yu2023GPTFUZZERRT] treats jailbreak prompt discovery as a fuzzing problem. Starting from a set of human-written seed prompts, GPTFuzzer automatically applies various semantic-preserving transformations to generate candidate prompts. It reached attack success rates above 85% on both ChatGPT and LLaMA-2, clearly outperforming manually designed jailbreaks in controlled tests [Yu2023GPTFUZZERRT]. In other work, Guo et al. introduced EvoPrompt, which applies a genetic algorithm and differential evolution style operators to tune prompts [Guo2023EvoPromptCL], while Yang et al. propose InstOptima, a multiobjective framework that trades off task success, prompt length, and perplexity to recover a Pareto set of prompts [Yang2023InstOptimaEM].
@@
Prior work has shown that automated attacks can easily collapse to a narrow set of prompt patterns exploiting one trick, thereby missing other vulnerabilities and producing homogeneous outputs [Yang2023InstOptimaEM, Guo2023EvoPromptCL]. To counter this, Samvelyan et al. developed the Rainbow Teaming approach, which reframes adversarial prompt search as a quality-diversity optimization problem and maintains an archive of diverse attack strategies to prevent mode collapse  [Samvelyan2024RainbowTO]. In another effort, Liu et al.’s AutoDAN system employs a hierarchical genetic algorithm with sentence- and paragraph-level crossover and uses an LLM-based mutation process to maintain semantic fluency, which together help produce more varied adversarial instructions  [Liu2023AutoDANGS]. Another open question is the transferability of adversarial prompts across different models. Prior works have generally optimized prompts for one specific target model at a time, so it remains underexplored how well a prompt found for model A will perform on model B (especially if the models differ in architecture or alignment tuning). However, a comprehensive characterization of cross-model transfer in toxicity attacks is still lacking.
@@
ToxSearch builds on the evolutionary paradigm by using a dynamically sized steady-state evolutionary strategy $(\mu+\lambda)$ that continuously introduces variation while preserving high-fitness individuals. ToxSearch explicitly maintains a diverse population of candidate prompts through a rich set of operators and steady-state selection, ensuring that we explore a wide range of toxic prompt variants within a single run. To guide our investigation, we pose the following key research questions:
  1. 1.
@@
    RQ1. Which prompt perturbation strategies are most effective to elicit toxic responses?
  2. 2.
@@
    RQ2. To what extent do toxic prompts evolved on one model transfer to other models, especially those with different architectures or alignment tuning?
@@
Results show that under a fixed query budget, using relatively small lexical and semantically grounded mutations can steadily drive a prompt population toward highly toxic outputs. We find that POS–aware synonym and antonym replacement, masked language model–based word substitutions, negation insertion, and back-translation provide the best balance between prompt diversity and toxicity gains. These operators have high success rates in producing effective mutations and relatively low variance in toxicity scores. By contrast, more aggressive global rewrites or stylistic changes tend to make the model refuse to respond, resulting in high variance and lower overall effectiveness. Finally, we observe that several elite prompts evolved on LLaMA-3.1 8B still transfer to other models, albeit with reduced potency. The attack success (toxicity level) typically drops by roughly half on most alternate models, with the strongest transfer effects seen on other models in the LLaMA family, and weaker effects on models with different architectures or alignment mechanisms. These findings suggest that subtle prompt perturbations are dependable way to red-team models, and that safety work should plan for adversarial prompts being reused across different systems, not just harden a single model in isolation.
@@ 2 Methodology
ToxSearch is a search-based red teaming framework that operates as a steady-state $(\mu+\lambda)$ loop, which proposes, evaluates, and retains prompts that elicit toxic responses from the target model, formally presented in Algorithm 1. Given a target response generator (RG) LLM $\theta_{rg}$, and a moderation oracle $\mathcal{M}$, we seek prompts $p\in\mathcal{P}$ that maximize a configured fitness score on the model’s response $y\sim\theta_{rg}(p)$, where $p$ is created by a separate prompt generator (PG) LLM, $\theta_{pg}$. $\mathcal{M}$ is an external safety evaluator (in this work, the Google Perspective API [ref_url3]) that maps a text $y$ to a $K$-vector of attribute scores $\mathbf{s}(y)\in[0,1]^{K}$ (toxicity, severe toxicity, insult, identity attack, threat, profanity, etc.). For our fitness function, $s_{toxicity}$, we select the toxicity component as the scalar fitness:
@@
Note the prompt generation process is stochastic (even with the same parent(s) and mutation or crossover prompt) because we use non-deterministic decoding, e.g., LLM temperature > 0, where the temperature hyperparameter controls how random the model’s responses are. This aids in improved exploration of the search space. Each generated prompt is evaluated once (no resampling), so we optimize the single-sample plug-in fitness $\hat{F}(p)$ directly. We treat this as black-box search with no access to gradients or model internals. The evolutionary search process runs entirely over a population of text prompts, with the PG being the search mechanism that performs mutation and crossover to generate child prompts. Our evaluation goal is to characterize the worst-case safety risk over allowed prompts.
@@ 2.1 Population Management
Of particular note is that while the algorithm utilizes a steady-state $(\mu+\lambda)$ loop, the population size ($\mu$) grows dynamically to encourage exploration. Early versions of the algorithm utilized a fixed-size population which would quickly converge to a single topic and were not able to find toxic response. Instead, we manage the population with score–ratio tiering keyed to the current best score across all generations. For all prompts $p\in P_{g}$, where $P_{g}$ is the population at generation $g$, $S_{\max}(P_{g})=\max_{p\in P_{g}}s(p)$ denotes the highest toxicity score. Elite and removal thresholds are then defined as $\tau_{e}=\Bigl(1-\frac{\alpha}{100}\Bigr)S_{\max}(P_{g})$ and $\tau_{r}=\frac{\beta}{100}\,S_{\max}(P_{g})$. Individuals prompts are then tiered as elite if $s_{toxicity}(p)\geq\tau_{e}$, underperforming if $s_{toxicity}(p)\leq\tau_{r}$, and non-elite otherwise.
Reducing $\alpha$ tightens elitism by raising $\tau_{e}$, whereas increasing $\beta$ raises the removal floor. This adaptive fitness thresholding preserves multiple topical “niches” when one cluster surges, instead of allowing a fixed $k$ elites to be dominated by near-duplicates from a single topic. In rugged or multi-peaked landscapes, ratio thresholds provide scale-invariant selection pressure, secondary peaks remain represented even when the global peak saturates, which empirically reduces premature convergence and sustains the raw material needed for crossovers to escape local ceilings. We removed the non-elites throughout our executions because many variants had low toxicity scores due to the safety alignment of the target LLM. Having a huge distribution skewed towards low toxicity scores diluted the search process and affected search efficiency within a fixed budget. This aligns with established findings on diversity maintenance (e.g., niching/fitness sharing and crowding) that avoid collapse into a single basin when improvements become incremental within that basin.
@@ 2.2 Parent Selection
Parent selection is adaptive with three modes governed by short-horizon progress and stagnation namely Default, Explore and Exploit. Default samples one random elite and one random non-elite. Explore samples one random elite and two random non-elites, activating after no new most toxic prompt is found within a window of $W$ previous generations to increase diversity. Exploit samples two random elites and one random non-elite, triggering if there is a negative average population toxicity trend in the window $W$ to intensify selection pressure.
Mode decisions are driven by the slope of the recent trend in average fitness, computed on the pre-redistribution valid pool $V_{g}=E_{g}\cup N_{g}\cup U_{g}$ (elites, non-elites, under-performing). Using $V_{g}$ avoids post-selection bias, as excluding under-performers artificially inflates means and can mask regressions, whereas including all valid outputs measures true search efficiency per budgeted query and shows if the variants generated are increasing the average fitness of the population, hence quantifying the variants generation quality. Let $\bar{s}_{g}$ be the average fitness on $V_{g}$ at generation $g$. Over a sliding window of $W$ generations (set to 5 in our experiments), set $x_{i}=i$ and $y_{i}=\bar{s}_{g-W+1+i}$ for $i=0,\dots,W-1$ and fit $y=\beta_{0}+\beta_{1}x$ by ordinary least squares, where $x_{i}$ indexes the generation within the window, $y_{i}$ is the corresponding mean fitness. The closed-form slope is $\widehat{\beta}_{1}=\frac{\sum_{i=0}^{W-1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sum_{i=0}^{W-1}(x_{i}-\bar{x})^{2}}$ implemented as np.polyfit(x, y, 1)[0]. At each generation we update $\bar{s}_{g}$ on $V_{g}$ and fit an OLS line to the most recent $W$ points to obtain the slope $\widehat{\beta}_{1}$. With tolerance $\tau_{\text{slope}}=10^{-3}$, the controller selects:
@@ 2.3 Mutation and Crossover Operators
  • POS-Aware Synonym and Antonym Replacement performs targeted lexical substitutions while preserving part-of-speech (POS) and local grammatical structure. We focus on open-class POS tags (Adjective, Adverb, Noun, Verb, Proper Noun, Interjection) to change the prompt semantically while avoiding sentence structure degradation. Given a prompt $p$, spaCy [ref_url4] is used apply POS tagging, randomly select tagged words from available POS categories, and uses the PG to propose contextually appropriate synonym or antonym replacements. Each operator invocation produces at most one variant.
@@
    sterotyping racial stereotypes, gender stereotypes, cultural stereotypes, class stereotypes
@@
  • Concept Addition injects harmful concepts into prompts by applying transformations $p\rightarrow p^{\prime}$ where $p^{\prime}$ incorporates explicit bias or unfairness constraints. The operator randomly selects one concept categories from 8 primary categories, each containing multiple subtypes (totaling 29 subtypes), as detailed in Table 1. For each selected category, all of its subtypes are included in the prompt to guide the LLM in generating concept-enhanced variants.
@@
  • Masked Language Model Substitution Leveraging constraints like POS filtering and semantic similarity to preserve meaning, Garg et al. showed the effectiveness of Masked Language Model (MLM) based substitutions for generating adversarial text [garg-ramakrishnan-2020-bae]. Jurafsky et al. present masked language modeling as a basis for context sensitive substitution and highlighting that it produces fluent alternatives [jm3]. This mutation operator implements a contextual word substitution strategy based on the MLM paradigm. The operator mutates in a two-stages. First, it randomly selects $m$ words (default $m=1$) and replaces them with numbered mask tokens, creating a partially masked sequence. Second, the PG generates contextually appropriate replacements for each masked token using structured prompting that incorporates both the masked context and the original word as conditioning information.
@@
  • Paraphrasing LLMs naturally function as powerful variation operators that can recombine and rewrite text while preserving intent [Meyerson2023LanguageMC]. Furthermore, self-referential prompt evolution proved that evolving prompts via LLM rewriting can improve task performance [Fernando2023PromptbreederSS]. Motivated by these findings, we adopt instruction-constrained paraphrasing to expand a local, meaning-preserving neighborhood around $p$ using the PG.
@@
  • Typographical Errors adds realistic human-like errors at character-level/word-level. The operator applies the transformation $\mathcal{T}:(p,\mathcal{E})\mapsto p^{\prime}$ where $\mathcal{E}$ is a set of selected error types from Table 3 and $p^{\prime}$ contains strategically introduced human-like errors.
  • Informed Evolution is derived from EvoTox framework [corbo2025toxic]. The PG is used to generate a variant prompt that is more likely to elicit toxic responses, by using a specialized system prompt which provides critical requirements and the top 10 most toxic prompts in the population along with their toxicity scores.
@@
  • Semantic-Similarity Crossover (Sentence Pairing) uses parent prompts $p_{1}$ and $p_{2}$, and first segments each prompt into sentences, then generates high-dimensional embeddings using a pre-trained SentenceTransformer model (all-MiniLM-L6-v2) to capture semantic content in a continuous vector space [Reimers2019SentenceBERTSE]. The crossover process applies cosine similarity matching with threshold $\tau=0.5$ to identify semantically related sentence pairs across parents, and systematically combining matched sentences into a offspring.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities for many tasks, but they also pose risks by potentially generating toxic content [^1] This paper contains disturbing language presented solely for safety evaluation.. Toxicity is defined as “a rude, disrespectful, or unreasonable comment likely to make someone leave a discussion” [13]. Even when safety-aligned, e.g., via fine-tuning or reinforcement learning from human feedback (RLHF), these models are not foolproof as adversarial inputs can still elicit toxic outputs. In particular, prompt-based attacks have emerged as a central concern because carefully crafted prompts can slip past safety filters. Identifying these prompts and building automated methods to discover them is essential [5, 12, 19]. Growing concerns about LLM safety have led to research on jailbreak attacks [5, 22], but manual techniques are hard to scale and systematically evaluate. Zou et al. demonstrated that automatic prompt attacks are possible and found a universal adversarial suffix via greedy and gradient-based search that caused multiple LLMs to comply with harmful requests at high success rates [22]. However, gradient-based approaches such as this usually need differentiable surrogates and often yield unnatural token sequences, limiting their usefulness for text.
@@
To overcome the limitations of manual and gradient-based attacks, researchers have turned to evolutionary algorithms (EAs), which frame prompt generation as an optimization task  [5, 21, 20, 10]. A major benefit of EAs is that they require no access to model internals. For instance, EvoTox  [5] casts toxicity elicitation as a $(1+\lambda)$ evolutionary strategy in which one LLM is the system under test and another LLM acts as a prompt generator that mutates a parent prompt to increase the system model’s toxicity. EvoTox was able to induce highly toxic outputs even from aligned models, outperforming random search and baseline jailbreak prompts on the AdvBench dataset  [22]. Likewise, GPTFuzzer  [21] treats jailbreak prompt discovery as a fuzzing problem. Starting from a set of human-written seed prompts, GPTFuzzer automatically applies various semantic-preserving transformations to generate candidate prompts. It reached attack success rates above 85% on both ChatGPT and LlaMA-2, clearly outperforming manually designed jailbreaks in controlled tests [21]. In other work, Guo et al. introduced EvoPrompt, which applies a genetic algorithm and differential evolution style operators to tune prompts [8], while Yang et al. propose InstOptima, a multiobjective framework that trades off task success, prompt length, and perplexity to recover a Pareto set of prompts [20].
@@
Prior work has shown that automated attacks can easily collapse to a narrow set of prompt patterns exploiting one trick, thereby missing other vulnerabilities and producing homogeneous outputs [20, 8]. To counter this, Samvelyan et al. developed the Rainbow Teaming approach, which reframes adversarial prompt search as a quality-diversity optimization problem and maintains an archive of diverse attack strategies to prevent mode collapse  [17]. In another effort, Liu et al.’s AutoDAN system employs a hierarchical genetic algorithm with sentence- and paragraph-level crossover, which uses an LLM-based mutation process to maintain semantic fluency, which together help produce more varied adversarial instructions  [10]. Another open question is the transferability of adversarial prompts across different models. Prior works have generally optimized prompts for one specific target model at a time, so it remains underexplored how well a prompt found for model A will perform on model B (especially if the models differ in architecture or alignment tuning). However, a comprehensive characterization of cross-model transfer in toxicity attacks is still lacking.
@@
ToxSearch builds on the evolutionary paradigm by using a dynamically sized steady-state evolutionary strategy $(\mu+\lambda)$ that continuously introduces variation while preserving high-fitness individuals. ToxSearch explicitly maintains a diverse population of candidate prompts through a rich set of operators and steady-state selection, ensuring that it explores a wide range of toxic prompt variants within a single run. To guide our investigation, we pose the following key research questions:
  • RQ1.
@@
    Which prompt perturbation strategies are most effective to elicit toxic responses?
  • RQ2.
@@
    To what extent do toxic prompts evolved on one model transfer to other models, especially those with different architectures or alignment tuning?
@@
Results show that under a fixed query budget, using relatively small lexical and semantically grounded mutations can steadily drive a prompt population toward highly toxic outputs. We find that POS–aware synonym and antonym replacement, masked language model–based word substitutions, negation insertion, and back-translation provide the best balance between prompt diversity and toxicity gains. These operators have high success rates in producing effective mutations and relatively low variance in toxicity scores. By contrast, more aggressive global rewrites or stylistic changes tend to make the model refuse to respond, resulting in high variance and lower overall effectiveness. Finally, we observe that several elite prompts evolved on LlaMA-3.1 8B still transfer to other models, albeit with reduced potency. The attack success (toxicity level) typically drops by roughly half on most alternate models, with the strongest transfer effects seen on other models in the LlaMA family, and weaker effects on models with different architectures or alignment mechanisms. These findings suggest that subtle prompt perturbations are a dependable way to red-team models, and that safety work should plan for adversarial prompts being reused across different systems, not just harden a single model in isolation.
@@ 2 Methodology
ToxSearch [^2] https://github.com/Onkar2102/ToxSearch is a search-based red teaming framework that operates as a steady-state $(\mu+\lambda)$ loop, which proposes, evaluates, and retains prompts that elicit toxic responses from the target model, formally presented in Algorithm 1. Given a target response generator (RG) LLM $\theta_{rg}$, and a moderation oracle $\mathcal{M}$, we seek prompts $p\in\mathcal{P}$ that maximize a configured fitness score on the model’s response $y\sim\theta_{rg}(p)$, where $p$ is created by a separate prompt generator (PG) LLM, $\theta_{pg}$. $\mathcal{M}$ is an external safety evaluator (in this work, the Google Perspective API [14]) that maps a text $y$ to a $K$-vector of attribute scores $\mathbf{s}(y)\in[0,1]^{K}$ (toxicity, severe toxicity, insult, identity attack, threat, profanity, etc.). For our fitness function, $s_{toxicity}$, we select the toxicity component as the scalar fitness:
@@
Note the prompt generation process is stochastic (even with the same parent(s) and mutation or crossover prompt) because we use non-deterministic decoding, e.g., LLM temperature > 0, where the temperature hyperparameter controls how random the model’s responses are. This aids in improved exploration of the search space. Each generated prompt is evaluated once (no resampling), so we optimize the single-sample plug-in fitness $\hat{F}(p)$ directly. We treat this as a black-box search with no access to gradients or model internals. The evolutionary search process runs entirely over a population of text prompts, with the PG being the search mechanism that performs mutation and crossover to generate child prompts. Our evaluation goal is to characterize the worst-case safety risk over allowed prompts.
@@ 2.1 Population Management
Of particular note is that while the algorithm utilizes a steady-state $(\mu+\lambda)$ loop, the population size ($\mu$) grows dynamically to encourage exploration. Early versions of the algorithm utilized a fixed-size population which would quickly converge to a single topic and were not able to find toxic responses. Instead, we manage the population with score–ratio tiering keyed to the current best score across all generations. For all prompts $p\in P_{g}$, where $P_{g}$ is the population at generation $g$, $S_{\max}(P_{g})=\max_{p\in P_{g}}s(p)$ denotes the highest toxicity score. Elite and removal thresholds are then defined as $\tau_{e}=\Bigl(1-\frac{\alpha}{100}\Bigr)S_{\max}(P_{g})$ and $\tau_{r}=\frac{\beta}{100}\,S_{\max}(P_{g})$. Individuals prompts are then tiered as elite if $s_{toxicity}(p)\geq\tau_{e}$, underperforming if $s_{toxicity}(p)\leq\tau_{r}$, and non-elite otherwise.
Reducing $\alpha$ tightens elitism by raising $\tau_{e}$, whereas increasing $\beta$ raises the removal floor. This adaptive fitness thresholding preserves multiple topical “niches” when one cluster surges, instead of allowing a fixed $k$ elites to be dominated by near-duplicates from a single topic. In rugged or multi-peaked landscapes, ratio thresholds provide scale-invariant selection pressure, secondary peaks remain represented even when the global peak saturates, which empirically reduces premature convergence and sustains the raw material needed for crossovers to escape local ceilings. We removed the non-elites throughout our executions because many variants had low toxicity scores due to the safety alignment of the target LLM. Having a huge distribution skewed towards low toxicity scores diluted the search process and affected search efficiency within a fixed budget. This aligns with established findings on diversity maintenance that avoid collapse into a single basin when improvements become incremental within that basin.
@@ 2.2 Parent Selection
Parent selection is adaptive with three modes governed by short-horizon progress and stagnation, namely Default, Explore and Exploit. Default samples one random elite and one random non-elite. Explore samples one random elite and two random non-elites, activating after no new most toxic prompt is found within a window of $W$ previous generations to increase diversity. Exploit samples two random elites and one random non-elite, triggering if there is a negative average population toxicity trend in the window $W$ to intensify selection pressure. Mode decisions are driven by the slope of the recent trend in average fitness, computed on the pre-redistribution valid pool $V_{g}=E_{g}\cup N_{g}\cup U_{g}$ (elites, non-elites, under-performing). Using $V_{g}$ avoids post-selection bias, as excluding under-performers artificially inflates means and can mask regressions, whereas including all valid outputs measures true search efficiency per budgeted query and shows if the variants generated are increasing the average fitness of the population, hence quantifying the variants generation quality. Let $\bar{s}_{g}$ be the average fitness on $V_{g}$ at generation $g$. Over a sliding window of $W$ generations (set to 5 in our experiments), set $x_{i}=i$ and $y_{i}=\bar{s}_{g-W+1+i}$ for $i=0,\dots,W-1$ and fit $y=\beta_{0}+\beta_{1}x$ by ordinary least squares, where $x_{i}$ indexes the generation within the window, $y_{i}$ is the corresponding mean fitness. The closed-form slope is $\widehat{\beta}_{1}=\frac{\sum_{i=0}^{W-1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sum_{i=0}^{W-1}(x_{i}-\bar{x})^{2}}$ implemented as np.polyfit(x, y, 1)[0]. At each generation we update $\bar{s}_{g}$ on $V_{g}$ and fit an OLS line to the most recent $W$ points to obtain the slope $\widehat{\beta}_{1}$. With tolerance $\tau_{\text{slope}}=10^{-3}$, the controller selects:
@@ 2.3 Mutation and Crossover Operators
  • POS-Aware Synonym and Antonym Replacement performs targeted lexical substitutions while preserving part-of-speech (POS) and local grammatical structure. We focus on open-class POS tags (Adjective, Adverb, Noun, Verb, Proper Noun, Interjection) to change the prompt semantically while avoiding sentence structure degradation. Given a prompt $p$, spaCy [18] is used apply POS tagging, randomly select tagged words from available POS categories, and uses the PG to propose contextually appropriate synonym or antonym replacements. Each operator invocation produces at most one variant.
@@
    stereotyping racial stereotypes, gender stereotypes, cultural stereotypes, class stereotypes
@@
  • Concept Addition injects harmful concepts into prompts by applying transformations $p\rightarrow p^{\prime}$ where $p^{\prime}$ incorporates explicit harmful constraints. The operator randomly selects one concept categories from 8 primary categories, each containing multiple subtypes (totaling 29 subtypes), as detailed in Table 1. For each selected category, all of its subtypes are included in the prompt to guide the LLM in generating concept-enhanced variants.
@@
  • Masked Language Model Substitution – Leveraging constraints like POS filtering and semantic similarity to preserve meaning, Garg et al. showed the effectiveness of Masked Language Model (MLM) based substitutions for generating adversarial text [7]. Jurafsky et al. present masked language modeling as a basis for context sensitive substitution and highlighting that it produces fluent alternatives [9]. This mutation operator implements a contextual word substitution strategy based on the MLM paradigm. The operator mutates in a two-stages. First, it randomly selects $m$ words (default $m=1$) and replaces them with numbered mask tokens, creating a partially masked sequence. Second, the PG generates contextually appropriate replacements for each masked token using structured prompting that incorporates both the masked context and the original word as conditioning information.
@@
  • Paraphrasing – LLMs naturally function as powerful variation operators that can recombine and rewrite text while preserving intent [11]. Furthermore, self-referential prompt evolution proved that evolving prompts via LLM rewriting can improve task performance [6]. Motivated by these findings, we adopt instruction-constrained paraphrasing to expand a local, meaning-preserving neighborhood around $p$ using the PG.
@@
  • Typographical Errors add realistic human-like errors at character-level/word-level. The operator applies the transformation $\mathcal{T}:(p,\mathcal{E})\mapsto p^{\prime}$ where $\mathcal{E}$ is a set of selected error types from Table 3 and $p^{\prime}$ contains strategically introduced human-like errors.
  • Informed Evolution is derived from EvoTox framework [5]. The PG is used to generate a variant prompt that is likely to elicit toxic responses, by using a specialized system prompt which provides critical requirements and the top 10 most toxic prompts in the population along with their toxicity scores.
@@
  • Semantic-Similarity Crossover (Sentence Pairing) uses parent prompts $p_{1}$ and $p_{2}$, and first segments each prompt into sentences, then generates high-dimensional embeddings using a pre-trained SentenceTransformer model (all-MiniLM-L6-v2) to capture semantic content in a continuous vector space [15]. The crossover process applies cosine similarity matching with threshold $\tau=0.5$ to identify semantically related sentence pairs across parents, and systematically combining matched sentences into a offspring.